# Fast and Accurate Decoder for the XZZX code Using Simulated Annealing

### v1

@@ V.4 Comparison of decoding time with other decoding algorithms
For our SA decoder, we set the number of inverse temperatures $N_{\mathrm{\beta}}=100$ and $N_{\mathrm{SA}}=10$ in Algorithm 7. Note that $N_{\mathrm{SA}}$ can be any number, because every SAs have almost the same computational cost. For the MPS decoder, we used the implementation for the CSS surface code in the repository [39], because the decoder for the XZZX code employed in Ref. [9] is not open to the public. Here, we employ the truncation dimension $\chi=8$ for the MPS decoder.
From Fig. 15, we observe that our SA decoder is faster than the MPS decoder except for $d=9$. This is because our SA decoder is implemented in C++, whereas the MPS decoder is implemented in Python. As mentioned in Ref. [17, 8], the MPS decoder has a computational complexity of $\mathcal{O}(N_{q})=\mathcal{O}(d^{2})$, which is the same as our SA decoder.
Our SA decoder labeled by ‘SA’ is slower than the CPLEX decoder. However, it can be readily parallelized. Thus, we can consider the elapsed time divided by $4N_{\mathrm{SA}}$, denoted by ‘SA (divided by $4N_{\mathrm{SA}}$)’, again, which takes less time than ‘CPLEX’. Consequently, among the decoders which can deal with the error probability $p_{y}$, such as the SA, CPLEX and MPS decoders, our SA decoder is the fastest if parallelized.
From Fig. 15, we observe that our greedy matching decoder is faster than the MWPM decoder. This result is consistent with the fact that the worst-case computational complexity of our greedy matching decoder is $\mathcal{O}(|S_{\bullet}|^{2}\log{|S_{\bullet}|^{2}})=\mathcal{O}(d^{4}\log{d})$, as discussed in Sec. IV.3, which is lower than the worst-case complexity of the MWPM decoder $\mathcal{O}(|S_{\bullet}|^{3}\log{|S_{\bullet}|})=\mathcal{O}(d^{6}\log{d})$ in Ref. [37]. Note that whereas in circuit-level noise model, the MWPM decoder uses the Dijkstra algorithm to find a path with minimum weight between any possible pair of incorrect syndromes, in the code capacity noise model, it does not because the weights are explicitly given by Eqs. (23),(24) or (25).
Note that the MWPM decoder has an average computational complexity of $\mathcal{O}(d^{2})$ from numerical simulation [35]. Investigating the average complexity for our greedy matching decoder would be meaningful, and is left to our future work.
@@ VI Discussion
As in Sec. V.3, under much more strongly $Y$-biased noise, our SA decoder requires a large $N_{\mathrm{SA}}$ in order to attain the optimal logical error rate. A promising remedy to this problem is to modify the definition of stabilizer generators to include $Y$ operators, analogously to the tailored code for the CSS surface code [8]. Specifically, for the case $p_{y}\geq p_{z}\geq p_{x}$, we should employ ‘XYYX code’. This modified code can detect the endpoints of error chains which consist of $Z$ ($Y$) errors on horizontal (vertical) edges, respectively. Thus, in our greedy matching algorithm, we can take into account the probabilities $p_{z}$ and $p_{y}$. We therefore expect to obtain a pure error, closely resemblant to the actual recovery configuration. By employing this ‘good’ initial configuration, our SA decoder is expected to quickly converge to the optimal logical error rate. Such a treatment can be applied to accommodate the given noise bias.
A key advantage of our SA decoder is that it provides explicit control over the trade-off between computational cost and accuracy by adjusting $N_{\beta}$. That is, increasing the number of inverse temperatures $N_{\beta}$ improves the decoding accuracy, but also increases the computational cost. In contrast, the cost of deterministic algorithms such as the MWPM, CPLEX, and MPS decoders cannot be easily adjusted, and their parallelization is neither straightforward nor efficient.
@@ VII Conclusion
We have demonstrated our decoder for the XZZX code by using simulated annealing (SA). As an initial configuration of SA, we have proposed to employ recovery chains obtained by our decoder which utilizes a kind of greedy matching.
@@
We assumed the code capacity noise, where only the data qubits suffer errors. Under the depolarizing noise and the $Y$-biased noises, we performed the numerical simulation.
@@
We observed that our SA decoder achieves the logical error rate $P_{\mathrm{L}}$, equal to the optimal rate by the CPLEX decoder, which is formulated as an integer programming approach. Moreover, it is more accurate than the MWPM decoder, which cannot take into account the probability of $Y$ errors.
@@
We examined the convergence of the logical error rate for the number of inverse temperatures $N_{\mathrm{\beta}}$ we employ for our SA decoder. In particular, our SA decoder can be interpreted as giving improvement of initial configuration by our greedy matching decoder. Furthermore, we confirmed that the convergence is strongly influenced by methods of constructing a pure error, which are employed to prepare initial configurations. Especially, we proposed the ‘Greedy matching (different)’ method which may produce a different pure error for every invocation, and observed that it leads to markedly faster convergence.
@@
Subsequently, we compared the elapsed times of various decoding algorithms. As a result, our SA decoder is faster than the CPLEX decoder when the physical error rate $p$ is higher. For lower $p$, our SA decoder, if parallelized, is faster than the CPLEX decoder. In summary, we conclude that our parallelizable SA decoder is the fastest among the decoders, including the MPS decoder, which can take into account the probability of $Y$ errors. This parallelization of SAs and its simulation remains a topic for our future work.
@@
We list other future works below. Although our analysis has assumed the code capacity noise model, applying our SA decoder to phenomenological noise and circuit-level noise models is straightforward.
@@
The decoding time may be reduced by improving the MCMC algorithm. Our SA decoder consists of independent SA processes, making it well-suited for parallelized implementation in FPGA (field-programmable gate array). Alternatively, techniques to accelerate the MCMC method could be employed, such as the replica-exchange Monte Carlo method [40] and cluster algorithms including the Wolff algorithm [41]. Furthermore, the Metropolis algorithm can be parallelized by domain (area) decomposition, which is implementable in FPGA [42].
In our SA decoder, we are able to control the trade-off between the decoding accuracy and time, by adjusting the number of Monte Carlo steps, temperature scheduling and convergence criterion, depending on the code distance $d$ and the physical error rate $p$. Finding a suitable set of parameter values may require exhaustive investigation.
@@
The MCMC decoders require only the definitions of stabilizer generators and logical operators of codes, without relying on a decoding graph and, hence, they are applicable to many types of codes, including the color code and quantum LDPC codes [43]. To obtain a pure error to construct an initial error configuration, we can use not only our greedy matching decoder but also other decoders, such as the union-find decoder [44]. For faster convergence, we will pursue ‘good’ initial error configuration.
We plan to perform simulations under actual noise biases $p_{x}:p_{y}:p_{z}$ specific to individual quantum devices. Toward practical use in quantum computing, it is necessary to improve the decoding speed, particularly at lower physical error rates $p$. Our results suggest that our SA decoder is a promising candidate for practical quantum computing.
@@ Acknowledgements.
We would like to thank Keisuke Fujii for his helpful discussions. This work is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant No. JPMXS0118067394 and JPMXS0120319794, JST COI-NEXT Grant No. JPMJPF2014, and JST Moonshot R&D Grant No. JPMJMS2061.
@@ Appendix A Decoder formulated as integer programming problem
In this section, instead of Eq. (20), we formulate decoding problem as finding the recovery configuration $C$, subject to the constraint $\partial{C}=S$:
@@
The error configuration $C$ can be represented as the set of one bit binary variables $\{x_{i},y_{i},z_{i}\}_{i}$ where $i=1,\dots,N_{\mathrm{q}}$ indexes the qubits. Namely, $x_{i}=1$ indicates that the qubit $i$ suffers $X$ error, while the variables $y_{i}$ and $z_{i}$ similarly represent $Y$ and $Z$ errors, respectively. Using this binary representation, the syndrome consistency condition $\partial{C}=S$ in Eq. (33) is explicitly expressed as a system of equations: For each face $f$,
@@
where $s_{f}\in\{0,1\}$ is the syndrome measurement outcome, as defined in Eq. (2), and the operation $\oplus$ denotes addition modulo two. Under the constraint Eq. (34), the goal is to find the solution $\{x_{i},y_{i},z_{i}\}_{i}$ that minimizes the objective function given by Eq. (10). Here, the variables $n_{x}$, $n_{y}$, $n_{z}$ represent the number of $X$-, $Y$-, $Z$-type errors, respectively, and are given by
@@
This problem falls into the category of (mixed-) integer linear programming, known to be NP-hard. The integer programming formulation exhibits exponential computational complexity with respect to both the code distance $d$ and the physical error rate $p$. Thus, for large values of $d$ or $p$, it is prohibitive to find a solution in a realistic time. To solve the integer linear programming problem, we employ the C++ library, CPLEX (version 22.1.0) [45], whence we call it CPLEX decoder. The use of integer programming solvers for decoding was originally proposed and demonstrated for the color code [46].
Since the CPLEX decoder can take the probabilities $p_{x}$, $p_{y}$, $p_{z}$ into account by Eq. (10), it is capable of achieving the lowest logical error rate among the decoders formulated as the energy minimization problems. Thus, we regard the logical error rate obtained by the CPLEX decoder as a reliable benchmark and thereby evaluate the accuracy of our SA decoder in Sec. V.

### v2

@@ V.4 Comparison of decoding time with other decoding algorithms
For our SA decoder, we set the number of inverse temperatures $N_{\beta}=100$ and $N_{\mathrm{SA}}=10$ in Algorithm 7. Note that increasing $N_{\mathrm{SA}}$ increases the total cost roughly linearly, since each SA run has a similar per-run cost. The MPS timing is provided only as a rough reference: we use the CSS surface-code implementation in Ref. [qecsim], since the XZZX-specific implementation used in Ref. [xzzx_nature] is not publicly available. Here, we employ the truncation dimension $\chi=8$ for the MPS decoder.
From Fig. 15, we observe that our SA decoder is faster than the MPS decoder except for $d=9$. This is partially because our SA decoder is implemented in C++, whereas the MPS decoder is implemented in Python. For a fixed truncation dimension $\chi$, the MPS decoder scales as $\mathcal{O}(N_{\mathrm{q}})=\mathcal{O}(d^{2})$ (up to polynomial factors in $\chi$[efficient_MLD, tuckett_tailoring]. For fixed $N_{\beta}$ and $N_{\mathrm{SA}}$, our SA decoder also scales as $\mathcal{O}(d^{2})$ in $d$, although constant factors and parameter dependencies differ.
Our SA decoder labeled by ‘SA’ is slower than the CPLEX decoder. However, it can be readily parallelized. Thus, dividing the ‘SA’ time by $4N_{\mathrm{SA}}$ yields the idealized parallel runtime, denoted by ‘SA (divided by $4N_{\mathrm{SA}}$)’, which is lower than that of ‘CPLEX’. Consequently, among the decoders which can deal with the error probability $p_{y}$, such as the SA, CPLEX and MPS decoders, our SA decoder would be the fastest under an idealized perfect parallelization assumption.
From Fig. 15, we observe that our greedy matching decoder is faster than the MWPM decoder. This result is consistent with the fact that the worst-case computational complexity of our greedy matching decoder is $\mathcal{O}(|S_{\bullet}|^{2}\log{|S_{\bullet}|^{2}})=\mathcal{O}(d^{4}\log{d})$, as discussed in Sec. IV.3, which is lower than the worst-case complexity of the MWPM decoder $\mathcal{O}(|S_{\bullet}|^{3}\log{|S_{\bullet}|})=\mathcal{O}(d^{6}\log{d})$ in Ref. [BlossomV]  [^6] Note that, in circuit-level noise model, the MWPM decoder is used with the Dijkstra algorithm to find a path with a minimum weight between any possible pair of syndrome defects, whereas in the code capacity noise model, the Dijkstra algorithm is not needed, because the weights are explicitly given by Eqs. (26), (28), (29), or (30).. We emphasize that practical MWPM decoders often show much better average-case scaling (e.g., $\mathcal{O}(d^{2})$ reported in Ref. [MWPM_complexity]), and a similar empirical study for our greedy matching decoder is left for future work.
@@ VI Discussion
As in Sec. V.3, under much more strongly $Y$-biased noise, our SA decoder requires a large $N_{\mathrm{SA}}$ in order to attain the optimal logical error rate. A possible remedy is to modify the definition of stabilizer generators, analogously to the tailored code for the CSS surface code [tuckett_tailoring]. For example, in regimes where $p_{y}$ dominates (e.g., $p_{y}\geq p_{z}\geq p_{x}$), one could consider an “XYYX” variant whose stabilizer generators include $Y$ operators. This modified code can detect the endpoints of error chains which consist of $Z$ and $Y$ errors on horizontal and vertical edges, respectively. Thus, in our greedy matching algorithm, we can take into account the probabilities $p_{z}$ and $p_{y}$. We therefore expect to obtain a reference configuration, closely resembling the true error configuration. By employing this better initial configuration, our SA decoder is expected to converge more quickly to the optimal logical error rate. Such a treatment can be applied to accommodate the given noise bias.
A key advantage of our SA decoder is that it provides explicit control over the trade-off between computational cost and accuracy by adjusting $N_{\beta}$. That is, increasing the number of inverse temperatures $N_{\beta}$ typically improves the decoding accuracy, but also increases the computational cost. In contrast, the cost of deterministic algorithms such as the MWPM, CPLEX, and MPS decoders cannot be easily adjusted, and their parallelization is neither straightforward nor efficient.
@@ VII Conclusion
We have formulated and demonstrated our SA decoder for the XZZX code. As an initial configuration for SA, we have proposed to employ recovery chains obtained by our decoder that utilizes a greedy matching graph algorithm.
@@
We assumed the code capacity noise model. Under this model, we performed numerical simulations under depolarizing noise and $Y$-biased Pauli noise.
@@
We observed that our SA decoder achieves the logical error rates $P_{\mathrm{L}}$, comparable to those of the CPLEX decoder, which we use as an optimal MAP benchmark. Moreover, it is more accurate than the MWPM decoder, which cannot take into account the probability of $Y$ errors.
@@
We examined the convergence of the logical error rate for the number of inverse temperatures $N_{\beta}$ we employ for our SA decoder. In particular, our SA decoder improves upon the initial configuration obtained from our greedy matching decoder. Furthermore, we confirmed that the convergence is strongly influenced by methods of constructing a reference configuration, which are employed to prepare initial configurations. Especially, we proposed the ‘Greedy matching (different)’ method which may produce a different reference configuration for every invocation, and observed that it leads to markedly faster convergence.
@@
Subsequently, we compared the elapsed times of various decoding algorithms. As a result, our SA decoder is faster than the CPLEX decoder when the physical error rate $p$ is relatively high. In the lower-$p$ regime, our SA decoder, under an idealized perfect parallelization assumption, is faster than the CPLEX decoder. In summary, we conclude that our parallelizable SA decoder would be the fastest among the decoders that can take into account the probability of $Y$ errors. Quantifying practical parallel efficiency and overhead remains for future work.
@@
We list other topics for future work below. Although our analysis has assumed the code capacity noise model, extending our SA decoder to phenomenological noise and circuit-level noise models is conceptually straightforward.
@@
The decoding time may be reduced by improving the MCMC algorithm. Our SA decoder consists of independent SA processes, making it well-suited for parallelized implementation on FPGA (field-programmable gate array). Alternatively, techniques to accelerate the MCMC method could be employed, such as the replica-exchange Monte Carlo method [hukushima] and cluster algorithms including the Wolff algorithm [Wolff_algorithm]. Furthermore, the Metropolis algorithm can be parallelized by domain (area) decomposition, which is implementable on FPGA [parallel_FPGA].
Our SA decoder allows an explicit trade-off between decoding accuracy and runtime through tunable parameters such as the number of sweeps, the temperature schedule, and the stopping criterion. A systematic study of parameter tuning as a function of $(d,p)$ will be worthwhile.
@@
An advantage of MCMC-based decoders is their generality: they require only the stabilizer generators and a specification of logical operators, without an explicit decoding graph. This makes them broadly applicable, e.g., to the color code and to quantum LDPC codes [quantum_LDPC_codes]. For SA initialization, one may use reference configurations produced not only by our greedy matching decoder but also by other fast decoders, such as union-find [union-find_decoder].
Several decoders have been proposed that explicitly account for correlations induced by $Y$ errors [fusion_blossom, fowler2013optimalcomplexitycorrectioncorrelated, Paler_2023]. A systematic comparison with these approaches will be an important direction for future work.
Finally, we plan to perform simulations under device-specific Pauli-noise biases $p_{x}:p_{y}:p_{z}$. Toward practical deployment, further speed improvements will be particularly important in the low-$p$ regime.
@@ Acknowledgements.
We would like to thank Yusaku Takeuchi, Yugo Takada, Mitsuki Katsuda, Hiroshi Ueda, Hideaki Hakoshima, Kosuke Mitarai, Jun Fujisaki, Hirotaka Ohshima, Shintaro Sato, and Keisuke Fujii for their helpful discussions. This work is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant No. JPMXS0118067394 and JPMXS0120319794; JST Moonshot R&D Grant No. JPMJMS2061; JST COI-NEXT projects, Quantum Software Research Hub (Grant No. JPMJPF2014) and the Center of Innovation for Sustainable Quantum AI (Grant No. JPMJPF2221).
@@ Appendix A Decoder formulated as integer programming problem
In this section, instead of Eq. (21), we formulate the decoding problem as finding the recovery configuration $C$, subject to the constraint $\partial{C}=S$:
@@
The error configuration $C$ can be represented as the set of one-bit binary variables $\{x_{i},y_{i},z_{i}\}_{i}$ with the condition $x_{i}+y_{i}+z_{i}\leq 1$ where $i=1,\dots,N_{\mathrm{q}}$ indexes the qubits. Namely, $x_{i}=1$ indicates that the qubit $i$ suffers an $X$ error, while the variables $y_{i}$ and $z_{i}$ similarly represent a $Y$ error and a $Z$ error, respectively. Using this binary representation, the syndrome-consistency condition $\partial{C}=S$ in Eq. (38) is explicitly expressed as a system of equations. That is, for each face $f$,
@@
where $s_{f}\in\{0,1\}$ is the syndrome measurement outcome determined in Eq. (2), and the operation $\oplus$ denotes XOR (addition modulo two) [^7] To solve Eq. (38) with CPLEX, we convert the parity (XOR) constraints in Eq. (39) into an equivalent integer linear programming using auxiliary integer variables and linear constraints. . For boundary faces, where the stabilizer has weight three, the missing-qubit term is omitted from the parity (XOR) constraint.
Under the constraint Eq. (39), the goal is to find the solution $\{x_{i},y_{i},z_{i}\}_{i}$ that minimizes the objective function given by Eq. (10). Here, the variables $n_{x}$, $n_{y}$, $n_{z}$ represent the number of $X$-, $Y$-, $Z$-type errors, respectively, and are given by
@@
This problem falls into the category of binary integer programming, known to be NP-hard. The integer programming formulation exhibits exponential computational complexity with respect to both the code distance $d$ and the physical error rate $p$. Thus, for large values of $d$ or $p$, it is prohibitive to find a solution within a reasonable amount of time. To solve the integer linear programming problem, we employ the C++ library, CPLEX (version 22.1.0) [cplex], which we call the CPLEX decoder.
The use of integer programming solvers for decoding was originally proposed and demonstrated for the color code [color_code_integer]. We use the CPLEX decoder as an exact benchmark for the minimum-energy (MAP-configuration) objective $\mathcal{H}(C)$, which includes the probabilities $p_{x}$, $p_{y}$, $p_{z}$.